3D TERRAIN RENDERING FOR GEOGRAPHIC INFORMATION SYSTEM]

Alvin Leung

Northeast Parallel Architectures Center October 1995

APPLICATION OF 3D TERRAIN RENDERING

3D terrain rendering for Geographic Information System  Application of 3D terrain rendering

Data – Landsat Image @ 30m + DEM model @ 100m Applications – Business, Education and Scientific visualization • Enjoy a high altitude fly-by of your city, state, country or Mars.

• Analyze business competition by showing the relation between population density, the location of competitors and competitor’s profit.

• Animation of weather formations, e.g. thunderstorms, on realistic back ground.

• Demonstrate the effect of geographic feature on weather and ocean current.

• Analyze the relationship between pollution and population density.

Alvin Leung, NPAC, alvin@npac.syr.edu  Oct 95, 3D-terrain-thepage